# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v12

@@ 4 Secure Training
Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
@@
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
@@
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
Table 1: Implementation Results for iDASH datasets with 10-fold CV
@@
(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
Table 2: Implementation Results for other datasets with 5-fold CV
@@ 5 Experiments
We implement the enhanced NAG based on HE with the library HEAAN. The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B . All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
@@
The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to (Kim et al., 2018a) for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for evaluation and encryption. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
6 Conclusion
@@
In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain. The proposed quadratic gradient is constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v13

@@ IV Secure Training
Both Adagrad and Adam algorithms are not practical solutions for homomorphic logistic regression due to their frequent reliance on inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
@@
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [3], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
@@
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. [3] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The initial weight vector $\beta^{(0)}$, initialized as a zero vector, and the diagonal elements of the matrix $\bar{B}$ are each replicated $n$ times to construct two corresponding matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
The public cloud receives the ciphertexts $\text{ct}Z$, $\text{ct}{\beta}^{(0)}$, and $\text{ct}{\bar{B}}$, and executes the enhanced NAG algorithm to iteratively refine the encrypted weight vector by updating $\text{ct}{\beta}^{(0)}$ toward a more optimal solution. Refer to [3] for a detailed description about how to calculate the gradient by HE programming.
@@
In the privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
TABLE I: Implementation Results for iDASH datasets with 10-fold CV
@@
[3] 5 7 0.04 6.07 62.87 0.689
TABLE II: Implementation Results for other datasets with 5-fold CV
@@ V Experiments
We implement the enhanced NAG algorithm over homomorphically encrypted data using the HEAAN library. The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B. All experiments on encrypted data were conducted on a public cloud instance equipped with 32 vCPUs and 64 GB of RAM.
For a fair comparison with the baseline in [3], we adopt the same 10-fold cross-validation (CV) technique on the iDASH dataset, which contains 1,579 samples with 18 features, and apply the same 5-fold CV setup to the remaining five datasets. Following [3], we report average accuracy and Area Under the Curve (AUC) as the primary evaluation metrics. The results of these experiments are summarized in Tables I and II, which also include average evaluation runtime and storage consumption (i.e., encrypted dataset for the baseline, and encrypted dataset plus $\bar{B}$ for our method). To ensure consistency in ciphertext storage, we adopt the same packing strategy proposed by Kim et al. [3], resulting in comparable ciphertext sizes, with the only addition being the ciphertexts required to encode $\bar{B}$. We use a learning rate of $1+0.9^{t}$ throughout all training procedures.
@@
The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32{,}768$, yielding a security level of $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to [3] for details on these parameters. Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm. Nevertheless, even with just four iterations, the proposed method achieves results that remain comparable to those of the baseline. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for encryption and evaluation. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
VI Conclusion
@@
In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain. The proposed quadratic gradient is preferably constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
As future work, we aim to establish a more rigorous theoretical analysis of the convergence properties of the enhanced algorithm. A more practical and application-driven direction involves systematically evaluating the performance of the enhanced method under different learning rate configurations.